OpenStudy (anonymous):

5). how do i fine the slope and y-intercept of the following? y=0, x=3 x=-1, y=0 Write the equation of the line in slope-intercept form.

OpenStudy (anonymous):

No y intercept. Slope is not defined

OpenStudy (anonymous):

y1 = 0 , y2 = 0 x1 = 3 and x2 =-1 (y2-y1) / (x2 - x1) = (0-0 ) / ( -1 -3 ) 0 / -4 ! not defined !!

OpenStudy (anonymous):

Is it y =0 x = 3 and y =-1, x =0 that way there will be a gradient and y intercept

OpenStudy (anonymous):

Then the gradient is 3 and the y intercept is (0,3) That makes the equation y = 3x + 3